OpenStudy (anonymous):

simplify by taking roots, if possible. sqrt324xy/3 sqrt3

OpenStudy (anonymous):

\[\sqrt{324xy}/3\sqrt{3}\]

OpenStudy (anonymous):

I think \[3\sqrt{4xy}\]

OpenStudy (anonymous):

??